# Eccentric features in the gravitational wave phase of dynamically formed black hole binaries

### v1

@@ 2.3 Numerical Procedure
Here we briefly lay out the numerical procedure we follow to here calculate the phase shift $d\phi$ in our setup. This model requires a total of 8 inputs: 4 from for the inner orbit BBH evolution ($M_{1},M_{2},a_{0},e_{0}$), and 4 for the outer orbit evolution of the BBH around the perturber ($M_{3},f_{m},a_{\mathrm{out}},e_{\mathrm{out}}$). First we derive the inner orbital evolution of the inspiralling BBH.
@@
  • We then compute the array for the corresponding merger time $t_{c}(e)$ given by Eq. 15.
@@
  • We choose orbital elements for the (non-decaying) outer orbit, $a_{\mathrm{out}}$ and $e_{\mathrm{out}}$, and an angular position for where the merger takes place, $f_{m}$.
  • The computed array of the BBH merger time, $t_{c}$, is then used in with $a_{\mathrm{out}}$, $e_{\mathrm{out}}$, and $f_{m}$, to numerically obtain an array for the corresponding true anomaly $f$.
@@
  • Finally, we now have matching arrays with $\Delta{t}$ and orbital time $T_{12}$, as a function of either time $t$, orbital eccentricity $e$, or true anomaly $f$. We then calculate the GW phase shift $d\phi$ given by Eq. 11.
@@ 2.4 Analytical Approximations
Before we move to the main results derived from our procedure outlined above, we here consider a few highly useful and general approximations to the GW phase shift that allow one to estimate its magnitude analytically and corresponding relevant scalings. We especially describe how these can be used for fast estimations of the phase shift for eccentric BBH mergers formed during chaotic few-body interactions, which are being explored both in controlled experiments (e.g. Samsing, 2018; Samsing et al., 2024) and in MC codes such as CMC (Kremer et al., 2020) where merger statistics are based on millions of scatterings.
Below we first present an analytical approximate solution to the general case of an inspiralling eccentric BBH moving on an arbitrary trajectory, while being subject to an arbitrary acceleration vector (see also Vijaykumar et al., 2023; Tiwari et al., 2024). We then consider solutions specific to inner and outer eccentric orbits applicable to PN N-body studies.
@@ 2.4.1 Phase Shift from a General Acceleration
We start by considering the case of a BBH moving on an arbitrary trajectory, as shown in Fig. 3. At the point of merger the BBH has a velocity $v_{m}$ and is subject to a general acceleration $\bf a$. While the velocity vector is tangential to the trajectory of the BBH, the acceleration vector does not have to be either parallel or perpendicular to the the trajectory. In a circular motion, the acceleration vector will be perpendicular to the velocity vector, but this is not the case for e.g. a simple eccentric orbit, where the velocity- and the acceleration vectors only are perpendicular to each other at peri- and apocentre. The question is what the maximum phase shift $d\phi$ can be for such a general trajectory and acceleration. We here estimate this to linear order, where we assume the acceleration and velocity is constant near the point of merger. As shown in Samsing et al. (2024), the maximum intrinsic value of the phase shift when the BBH inspirals with non-zero eccentricity, e.g. without taking into account the observer position, will arise close to merger, which justifies this assumption.
Figure 3: Schematic depicting a binary with an accelerated centre of mass (turquoise line) with respect to an inertial reference binary (orange line). The presence of curvature in the binary’s path creates a time varying Doppler shift that influences its orbital frequency, as measured by an observer along a given line of sight (dotted red line).
@@ 2.4.2 Circular Approximation
The approximation above is at the linear level. However, when considering the specific case of a BBH moving on an eccentric orbit it is possible to provide a slightly more accurate, but still fast, estimate by approximating the trajectory by a circle around the point of merger instead of a line. For this, we continue under the assumption that the BBH is moving on circle that is tangent to the point of merger, with radius $r_{m}$, where $r_{m}$ is the distance between the BBH COM and the third perturbing object, and velocity $v$ defined as,
$v=\sqrt{GM/r_{m}}.$ (20)
@@
$\theta(t)=\frac{vt}{r_{m}},$ (21)
the RD is found using standard geometry to be,
$\begin{split}\Delta t(t)&=2\mu R[(\theta/2)^{2}+\sin^{2}(\theta/2)\\ &\qquad-2(\theta/2)\sin(\theta/2)\cos(\theta/2)]^{1/2}.\end{split}$ (22)
@@
$\begin{split}d\phi&\approx 2\frac{\sqrt{Gm_{12}}\mu r_{m}}{a(e)^{3/2}}\bigg{[}\frac{GM}{4r_{m}^{3}}t_{c}(e)^{2}+\sin^{2}\left(\sqrt{\frac{GM}{r_{m}}}\frac{t_{c}(e)}{2r_{m}}\right)\\ &\qquad-\sqrt{\frac{GM}{r_{m}}}\frac{t_{c}(e)}{r_{m}}\sin\left(\sqrt{\frac{GM}{r_{m}}}\frac{t_{c}(e)}{2r_{m}}\right)\\ &\qquad\quad\times\cos\left(\sqrt{\frac{GM}{r_{m}}}\frac{t_{c}(e)}{r_{m}}\right)\bigg{]}^{1/2},\end{split}$ (23)
where $t_{c}$ is given by Eq. 15. If one uses the eccentricity as evolution factor, one can derive $d\phi(e)$ in closed form for any value of $e$ using the above equation, together with Eq. 13 and Eq. 15. This will be shown later in Sec. 4.
@@ 2.4.3 Fast Linear Approximation
The input $r_{0}$ is well defined and easy to estimate, in contrast to $a_{0}$ and $e_{0}$. Combining these relations, we reach the following relation (Samsing et al., 2024),
@@ 2.5 Limitations and Conclusions
Both the linear and the circular approximations from above are based on assumptions of constant velocity and acceleration during the time over which the phase shift is observed, which are generally justified. However, in some cases, especially relevant for the LISA mission, the observational time window can be comparable or even longer than the outer orbital time, which leads to periodic features from the time-evolving Doppler shift that will naturally enter the GW waveform and phase shift (Samsing et al., 2024). In our considered case, where dynamically assembled binaries are often evolving along highly eccentric outer orbits, other effects can occur on much smaller timescales, as the acceleration will vary greatly over the orbit from pericentre to apocentre. Most notably, if the BBH happens to pass near pericentre at the time of observation, one expects large non-linear dephasing effects on a timescale set by the time of pericentre passage, which can be orders-of-magnitude smaller than the outer orbital time. Such effects are not captured by our approximations from this section, but below we will study such encounters in much more detail for the first time.
@@ 3.1 Imprints of Orbital Eccentricity
Figure 4: Top 4 panels: trajectories of four different phase shift-inducing scenarios, each with different $f_{m}$. The eccentricity of the outer orbit is 0.6. The turquoise curve is the true trajectory of the binary (merging at the smaller black dot at the corresponding $f_{m}$) on an eccentric orbit around BH3 (the larger black dot). The reference trajectory in isolation is depicted by the straight orange line. The dashed straight lines represent the distance $l(t)$ (Eq. 10) between the two trajectories which is used to calculate the time delay according to $\Delta t(t)=l(t)/c$. The red cross corresponds to the place on the orbit where $d\phi_{\mathrm{max}}$ occurs, which is quite close to merger in all cases. We also show in pink the circle that has $R=r_{m}$, which we use for our circular approximation. The black and grey dashed ellipses depict the Keplerian orbits of the binary COM and BH3, respectively. Bottom panel: the phase shift $d\phi$ as a function of time for each scenario. The black dashed lines in this panel represent the temporal location of the same lines in the trajectory plots.
Figure 5: Top 4 panels: trajectories of four different phase shift-inducing scenarios, each with different $f_{m}$. The eccentricity of the outer orbit is 0.9. The turquoise curve is the true trajectory of the binary (merging at the smaller black dot at the corresponding $f_{m}$) on an eccentric orbit around BH3 (the larger black dot). The reference trajectory in isolation is depicted by the straight orange line. The dashed straight lines represent the distance $l(t)$ (Eq. 10) between the two trajectories which is used to calculate the time delay according to $\Delta t(t)=l(t)/c$. The red cross corresponds to the place on the orbit where $d\phi_{\mathrm{max}}$ occurs, which is quite close to merger in all cases. We also show in pink the circle that has $R=r_{m}$, which we use for our circular approximation. The black and grey dashed ellipses depict the Keplerian orbits of the binary COM and BH3, respectively. Bottom panel: the phase shift $d\phi$ as a function of time for each scenario. The black dashed lines in this panel represent the temporal location of the same lines in the trajectory plots.
@@
$\frac{d{\phi}_{\mathrm{apo}}}{d{\phi}_{\mathrm{peri}}}\approx(1+e)^{2}/(1-e)^{2},$ (33)
and correspondingly by a factor of $1/(1-e)^{2}$ compared to if the BBH was circular with $R=a_{\mathrm{out}}$.
@@
Figure 6: Top: In grey, we show the dependence of $d\phi_{\mathrm{max}}$ on the true anomaly at merger $f_{m}$. The pink curve represents our circular approximation (Eq. 23). The outer eccentricity is 0.6. Bottom: the same relation, this time with $e_{\mathrm{out}}=0.9$.
In summary, the leading order effect from introducing a small to moderate eccentricity of the outer orbit, is a change in the maximum phase shift $d\phi_{\mathrm{max}}$. Generally, as the distance $r$ between the binary and third object at the point of merger increases, the peak value decreases as $\propto 1/r^{2}$. Note here that there will be further dynamically constraints on the BBH if the outer orbit eccentricity increases. We will touch upon this in Sec. 4.2, but a more throughout discussion will be saved for future numerical studies.
@@ 3.2 Strong Effects and Secondary Peaks
We focus here on the new qualitative features that go beyond the expectations from the $1/r_{m}^{2}$ scaling. For highly eccentric outer orbits, the maximum phase shift is achieved for binaries that merge shortly after their pericentre passage. As an example, a binary merging at $f_{m}\sim-0.2\pi$ passes the high curvature region at pericentre, changing direction rapidly. This change in direction, combined with a high orbital velocity, causes higher possible phase shifts as opposed to a binary simply merging at pericentre ($f_{m}=0$). In the regions $-0.75\pi<f_{m}<-0.25\pi$ and $0.25\pi<f_{m}<0.75\pi$, the circular approximation respectively under- and overestimates the maximum phase shift due to the extra ac- and deceleration of the binary from the strong curvature.
Figure 7: Schematic visualisation of the secondary peak. In orange, we depict the trajectory of a binary merging before pericentre, giving rise to the typical single-peak phase shift. The trajectory of a binary merging after pericentre is shown in pink, where the pericentre passage gives rise to a second peak.
@@
More characteristics unique to highly eccentric outer orbits are visible in Fig. 5. The case $f_{m}=-0.75\pi$ has two peaks: a main peak that occurs at around the same time as for the other curves, and a secondary peak at around $6000$ seconds before merger. In Fig. 7, we schematically depict this scenario. In orange, we show the trajectory of a binary that produces a single-peak phase shift. The case with the secondary peak is shown in pink. The position of merger on the orbit of this binary is such, that the binary passes pericentre right before the main peak close to merger. The strong curvature effects from the pericentre passage lead to a boost in time delay, resulting in a second peak in the phase shift. There is an $f_{m}$ sweetspot on the outer orbit where the secondary peak may actually be higher than the main one. For this to happen, $f_{m}$ needs to be far enough from pericentre that the pericentre passage happens before the main peak, but close enough to pericentre that the distance is not to large. In our considered setup, the sweet spot is at $f_{m}\sim-0.2\pi$. We stress that this is an important feature of a system like this, as it is a direct mapping of the environment in which the BBH mergers. This only occurs when $e_{\mathrm{out}}$ is high and only occurs in a specific part of the outer orbit so a possible observation of this would undoubtedly pin down properties of this BBH formation channel.
@@
In Fig. 8 we show in more detail how the phase shift curves vary around pericentre with changing $f_{m}$. In the top panel, we see that curves with the highest peak in the $e_{\mathrm{out}}=0.6$ scenario are green/light blue, which correspond to mergers at pericentre. The colour of the highest peak at $e_{\mathrm{out}}=0.9$ is a deeper blue, corresponding to mergers just after pericentre (at $f_{m}\sim-0.2\pi$). The bottom panel visualises how the secondary peak evolves. As we move from pink into blue, the merger happens increasingly close to pericentre, so the secondary peak slowly melts together with the main peak.
@@
Fig. 9 visualises the shape of $d\phi$, including the changing location of the maximum, for a full range of $f_{m}$ at a high outer eccentricity. Each vertical line in this plot represents the path of a binary merging at a specific $f_{m}$; the binary forms at the top, and eventually merges at the bottom. We show the range of true anomaly $f$ as a function of $f_{m}$. The colour represents the (normalised) phase shift at each $f$ in the orbit. There are clear variations in the location on the outer orbit where the maximum $d\phi$ occurs. It is furthest away from merger at around $f_{m}\sim 0.25\pi$. It is this region where the acceleration on the orbit is highest so it covers the largest range of $f$ between peak and merger. It is closest to merger around apocentre, as the orbital speed is so low that it barely moves on the orbit. Between $f_{m}\sim-0.75\pi$ and $f_{m}\sim-0.50\pi$ the presence of the secondary peak is clearly visible. We see an imprint of the pericentre passage with the secondary peak occurring just before pericentre, at $f_{m}\sim 0.2\pi$. The secondary and primary peak gradually merge when we move to the right from $f_{m}=-0.75\pi$.
@@
Figure 8: The phase shift $d\phi$ as a function of time, for a range of $f_{m}$ close to pericentre ($-0.75\pi\leq f_{m}\leq 0.75\pi$). The ellipse in the top left corner illustrates the outer eccentric orbit. Every coloured line represent a certain $f_{m}$, where the location of merger is the point where that line intersects with the ellipse. The colour of each line in the main plot corresponds to the colour of the lines within the ellipse. Top: We show the results for $e_{\mathrm{out}}=0.6$. Bottom: Results for $e_{\mathrm{out}}=0.9$.
Figure 9: Heat map of the the phase shift, as a function of the orbital phase $f$, for different realisations of $f_{m}$. Each vertical line represents the trajectory of an inspiralling binary, forming at some true anomaly $f_{f}$ (top), and merging at $f_{m}$ (bottom). We normalise $d\phi$ for each trajectory, such that $d\phi_{\mathrm{max}}=1$ for every $f_{m}$. We employ an outer eccentricity of 0.9.

### v2

@@ 2.3 Numerical Procedure
Figure 3: Binary undergoing accelerated motion. Schematic depicting a binary with an accelerated centre of mass (turquoise line) with respect to an inertial reference binary (orange line). The presence of curvature in the binary’s path creates a time varying Doppler shift that influences its orbital frequency, as measured by an observer along a given line of sight (dotted red line).
Here we briefly lay out the numerical procedure we follow to calculate the phase shift $d\phi$ in our setup. Our model requires a total of 8 inputs: 4 from for the inner orbit BBH evolution ($m_{1},m_{2},a_{0},e_{0}$), where $a_{0},e_{0}$ here refer to the initial values right after assembly, and 4 for the outer orbit evolution of the BBH around the perturber ($m_{3},f_{m},a_{\mathrm{out}},e_{\mathrm{out}}$). First we derive the inner orbital evolution of the inspiralling BBH.
@@
  • We then compute the array for the corresponding merger time $t_{c}$ given by Eq. 14.
@@
  • We first choose orbital elements for the (non-decaying) outer orbit, $a_{\mathrm{out}}$ and $e_{\mathrm{out}}$, and an angular position for where the merger takes place, $f_{m}$.
  • The computed array of the BBH merger time, $t_{c}$ (from the inner binary as described above) is then used with $a_{\mathrm{out}}$, $e_{\mathrm{out}}$, and $f_{m}$, to numerically obtain an array for the corresponding true anomaly $f$.
@@
  • Finally, we now have matching arrays with $\Delta{t}$ and orbital time $T_{12}$, as a function of either time $t$, orbital eccentricity $e$, or true anomaly $f$. We then calculate the maximum GW phase shift $d\phi$ given by Eq. 10.
@@ 2.4 Analytical Approximations
Before we move to the main results derived from our procedure outlined above, we first consider a few highly useful and general approximations to the GW phase shift that allow one to estimate its magnitude analytically and corresponding relevant scalings. We especially describe how these can be used for fast estimations of the phase shift for eccentric BBH mergers formed during chaotic few-body interactions, which are being explored both in controlled experiments (e.g. Samsing, 2018; Samsing et al., 2024) and in MC codes such as CMC (Kremer et al., 2020) where merger statistics are based on millions of scatterings.
Below we first present an analytical approximate solution to the general case of an inspiralling eccentric BBH moving on an arbitrary trajectory, while being subject to an arbitrary acceleration vector (for recent work on circular binaries, see also Vijaykumar et al., 2023; Tiwari et al., 2024b). We then consider solutions specific to inner and outer eccentric orbits applicable to PN N-body studies.
@@ 2.4.1 Phase Shift from a General Acceleration
We start by considering the case of a BBH moving on an arbitrary trajectory, as shown in Fig. 3. At the point of merger the BBH has a velocity $v_{m}$ and is subject to a general acceleration $\bf a$. While the velocity vector is tangential to the trajectory of the BBH, the acceleration vector does not have to be either parallel or perpendicular to the the trajectory. In a circular motion, the acceleration vector will be perpendicular to the velocity vector, but this is not the case for e.g. a simple eccentric orbit, where the velocity- and the acceleration vectors only are perpendicular to each other at peri- and apocentre. The question is what the maximum phase shift $d\phi$ can be for such a general trajectory and acceleration (see also Tiwari et al. (2024a)). We here estimate this to linear order, where we assume the acceleration and velocity is constant near the point of merger. As shown in Samsing et al. (2024), the maximum value of the phase shift when the BBH inspirals with non-zero inner orbit eccentricity, e.g. without taking into account the observer position, will arise close to merger, which justifies this assumption.
@@ 2.4.2 Circular Approximation
The approximation above is at the linear level. However, when considering the specific case of a BBH moving on an eccentric orbit it is possible to provide a slightly more accurate, but still fast, estimate by approximating the trajectory by a circle around the point of merger instead of a line. For this, we continue under the assumption that the BBH is moving on a circle that is tangent to the point of merger with a radius corresponding to the distance between the BBH and the three-body COM, $\mu r_{m}$, and effective velocity $v$ defined as,
$v=\sqrt{Gm_{123}/r_{m}}.$ (19)
@@
$\theta(t)={vt}/{r_{m}},$ (20)
the RD is found in this approximation using standard geometry (see Samsing et al. (2024) for further details),
$\begin{split}\Delta t(t)&=\frac{2\mu r_{m}}{c}[(\theta/2)^{2}+\sin^{2}(\theta/2)\\ &\qquad-2(\theta/2)\sin(\theta/2)\cos(\theta/2)]^{1/2}.\end{split}$ (21)
@@
$\begin{split}d\phi&\approx 2\frac{\sqrt{Gm_{12}}\mu r_{m}}{a(e)^{3/2}}\bigg{[}\frac{Gm_{123}}{4r_{m}^{3}}t_{c}(e)^{2}+\sin^{2}\left(\sqrt{\frac{Gm_{123}}{r_{m}}}\frac{t_{c}(e)}{2r_{m}}\right)\\ &\qquad-\sqrt{\frac{Gm_{123}}{r_{m}}}\frac{t_{c}(e)}{r_{m}}\sin\left(\sqrt{\frac{Gm_{123}}{r_{m}}}\frac{t_{c}(e)}{2r_{m}}\right)\\ &\qquad\quad\times\cos\left(\sqrt{\frac{Gm_{123}}{r_{m}}}\frac{t_{c}(e)}{r_{m}}\right)\bigg{]}^{1/2},\end{split}$ (22)
where $t_{c}$ is given by Eq. 14. If one uses the eccentricity as evolution factor, one can derive $d\phi(e)$ in closed form for any value of $e$ as described in the following.
Figure 4: GW phase shifts from eccentric orbits. Top 4 panels: Trajectories of four different GW phase shift-inducing scenarios, each with different $f_{m}$. The eccentricity of the outer orbit is set to $e_{\rm out}=0.6$. The turquoise curve is the true trajectory of the binary (merging at the smaller black dot at the corresponding $f_{m}$) on an eccentric orbit around BH3 (the larger black dot). The reference trajectory in isolation is depicted by the straight orange line. The dashed straight lines represent the distance $l(t)$ between the two trajectories which is used to calculate the time delay according to $\Delta t(t)=l(t)/c$. The red cross corresponds to the place on the orbit where $d\phi_{\mathrm{max}}$ occurs, which is quite close to merger in all cases. We also show in pink the circle with radius equal to $\mu r_{m}$, which we use for our circular approximation. The black and grey dashed ellipses depict the Keplerian orbits of the binary COM and BH3, respectively. Bottom panel: the phase shift $d\phi$ as a function of time for each scenario. The black dashed lines in this panel represent the temporal location of the same lines in the trajectory plots.
Figure 5: GW phase shifts from eccentric orbits. Similar to Fig. 4, but with the eccentricity of the outer orbit set to $e_{\rm out}=0.9$.
@@ 2.4.3 Fast Linear Approximation
The input $r_{0}$ is well defined and easy to estimate, in contrast to $a_{0}$ and $e_{0}$. Combining these relations, we reach the following relation,
@@ 2.5 Limitations and Conclusions
Both the linear and the circular approximations from above are based on assumptions of constant velocity and acceleration during the time over which the phase shift is observed, which are generally justified. However, in some cases, especially relevant for the LISA mission, the observational time window can be comparable or even longer than the outer orbital time, which leads to periodic features from the time-evolving Doppler shift that will naturally enter the GW waveform and phase shift. In our considered case, where dynamically assembled binaries are often evolving along highly eccentric outer orbits, other effects can occur on much smaller timescales, as the acceleration will vary greatly over the orbit from pericentre to apocentre. Most notably, if the BBH happens to pass near pericentre at the time of observation, one expects large non-linear dephasing effects on a timescale set by the time of pericentre passage, which can be orders-of-magnitude smaller than the outer orbital time. Such effects are not captured by our approximations from this section, but below we will study such encounters in much more detail.
@@ 3.1 Imprints of Orbital Eccentricity
$\frac{d{\phi}_{\mathrm{peri}}}{d{\phi}_{\mathrm{apo}}}\approx\frac{(1+e)^{2}}{(1-e)^{2}},$ (32)
and correspondingly by a factor of $1/(1-e)^{2}$ compared to if the BBH was circular with radius $a_{\mathrm{out}}$.
@@
Figure 6: Maximum GW phase shift. Top: In grey, we show the dependence of $d\phi_{\mathrm{max}}$ on the true anomaly at merger $f_{m}$. The pink curve represents our circular approximation (Eq. 22). The outer eccentricity is 0.6. Bottom: the same relation, this time with $e_{\mathrm{out}}=0.9$.
In summary, the leading order effect from introducing a small to moderate eccentricity of the outer orbit, is a change in the maximum phase shift $d\phi_{\mathrm{max}}$. Generally, as the distance $r$ between the binary and third object at the point of merger increases, the peak value decreases as $\propto 1/r^{2}$. Note here that there will be further dynamically constraints on the BBH if the outer orbit eccentricity increases. We will touch upon this in Sec. 4.2.
@@ 3.2 Strong Effects and Secondary Peaks
We focus here on the new qualitative features that go beyond the expectations from the $1/r_{m}^{2}$ scaling. For highly eccentric outer orbits, the maximum phase shift is achieved for binaries that merge shortly after their pericentre passage. As an example, a binary merging at $f_{m}\sim-0.2\pi$ passes the high curvature region at pericentre, changing direction rapidly. This change in direction, combined with a high orbital velocity, causes higher possible phase shifts as opposed to a binary simply merging at pericentre ($f_{m}=0$). In the regions $-0.75\pi<f_{m}<-0.25\pi$ and $0.25\pi<f_{m}<0.75\pi$, the circular approximation respectively under- and overestimates the maximum phase shift due to the extra ac- and deceleration of the binary from the strong curvature. This is futher outlined in Fig. 6.
Figure 7: GW phase shift, eccentricity, and double peaks. The phase shift $d\phi$ as a function of time, for a range of $f_{m}$ close to pericentre ($-0.75\pi\leq f_{m}\leq 0.75\pi$). The ellipse in the top left corner illustrates the outer eccentric orbit. Every coloured line represent a certain $f_{m}$, where the location of merger is the point where that line intersects with the ellipse. The colour of each line in the main plot corresponds to the colour of the lines within the ellipse. Top: We show the results for $e_{\mathrm{out}}=0.6$. Bottom: Results for $e_{\mathrm{out}}=0.9$.
@@
More characteristics unique to highly eccentric outer orbits are visible in Fig. 5. The case $f_{m}=-0.75\pi$ has two peaks: a main peak that occurs at around the same time as for the other curves, and a secondary peak appearing at earlier times. In Fig. 9 we schematically depict this scenario, where in orange we show the trajectory of a binary that produces a single-peak phase shift, and in pink a case where the merger happens right after peri-center passage, which gives rise to the earlier second peak. There is an $f_{m}$ sweetspot on the outer orbit where the secondary peak may actually be higher than the main one. For this to happen, $f_{m}$ needs to be far enough from pericentre that the pericentre passage happens before the main peak, but close enough to pericentre that the distance is not to large. In our considered setup, the sweet spot is at $f_{m}\sim-0.2\pi$. We stress that this is an important feature of a system like this, as it is a direct mapping of the environment in which the BBH merges. This only occurs when $e_{\mathrm{out}}$ is high and only occurs in a specific part of the outer orbit so a possible observation of this would undoubtedly pin down properties of this BBH formation channel.
@@
In Fig. 7 we show in more detail how the GW phase shift curves vary around pericentre with changing $f_{m}$, where each colored line corresponds to a different value of $f_{m}$, as illustrated in the upper left insert figure for each panel ($f_{m}$ for a given coloured line is where the line ends at the outer orbit shown with a black dashed line). In the top panel, we see that curves with the highest peak in the $e_{\mathrm{out}}=0.6$ scenario are green/light blue, which correspond to mergers near pericentre. The colour of the highest peak at $e_{\mathrm{out}}=0.9$ is a deeper blue, corresponding to mergers just after pericentre (at $f_{m}\sim-0.2\pi$). The bottom panel visualises how the secondary peak evolves. As we move from pink into blue, the merger happens increasingly close to pericentre, so the secondary peak slowly melts together with the main peak.
@@
Fig. 8 visualises the shape of $d\phi$, including the changing location of the maximum, for a full range of $f_{m}$ at a high outer eccentricity. Each vertical line in this plot represents the GW phase shift path of a binary merging at a specific $f_{m}$, where the binary evolves as a function of true anomaly $f$ from the top and down. The colour represents the (normalised) phase shift at each $f$ in the orbit. It is clearly seen how the evolving GW phase shift depends on how and where the BBH merges on the outer eccentric orbit. In particular, between $f_{m}\sim-0.75\pi$ and $f_{m}\sim-0.50\pi$ the presence of the secondary peak is clearly visible.
@@
Figure 8: Orbital phase and GW phase shift. Heat map of the the phase shift, as a function of the orbital phase $f$, for different realisations of $f_{m}$. Each vertical line represents the trajectory of an inspiralling binary, forming at some true anomaly $f_{f}$ (top), and merging at $f_{m}$ (bottom). We normalise $d\phi$ for each trajectory, such that $d\phi_{\mathrm{max}}=1$ for every $f_{m}$. We employ an outer eccentricity of 0.9.
Figure 9: Formation of double peak mergers. Schematic visualisation of the secondary peak. In orange, we depict the trajectory of a binary merging before pericentre, giving rise to the typical single-peak phase shift. The trajectory of a binary merging after pericentre is shown in pink, where the pericentre passage gives rise to a second peak.